OpenStudy (anonymous):

rewrite this formula to solve for the percent sales tax t=c+cb t=total price of the shirt c=cost of shirt b=sales tax

OpenStudy (perl):

b = (t - c ) /c

OpenStudy (igreen):

t = c + cb You want to get 'b' by itself. So first you subtract 'c' to both sides: t - c = cb Now divide c to both sides, what do you get?

OpenStudy (igreen):

@jancilde

OpenStudy (igreen):

@perl Don't give out answers.

OpenStudy (anonymous):

t-c/c=b

OpenStudy (anonymous):

^ rebel xD

OpenStudy (perl):

@iGreen check out my profile

OpenStudy (igreen):

Yep, you got it. @jancilde

OpenStudy (igreen):

@perl What you're doing is against the CoC.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @jancilde

OpenStudy (perl):

@iGreen you gave the answer too

OpenStudy (perl):

well, except for the last step . you basically handed and delivered the answer

OpenStudy (igreen):

So? I can still do that it's not giving them the answer. You're just giving the answer and you don't even have an explanation.

OpenStudy (patrickweewoos):

@perl i need help with 2 more